OpenStudy (anonymous):

Integrate (e^3x + 2e^2x)/(e^3x + 3e^2x + 9) dx from 0 to 1.

OpenStudy (anonymous):

\[\int_0^1\frac{e^{3x}+2e^{2x}}{e^{3x}+3e^{2x}+9}~dx\] Let \(u=e^{3x}+3e^{2x}+9\), then \(du=\left(3e^{3x}+6e^{2x}\right)~dx\) \[\frac{1}{3}\int_0^1\frac{3e^{3x}+6e^{2x}}{e^{3x}+3e^{2x}+9}~dx={\large\frac{1}{3}\int_{13}^{e^3+3e^2+9}\frac{du}{u}}\]